Mr. Karim is planning to pursue MBA studies. Initial inquiry shows that he will have to pay RM 24,000 as tuition fee, buy books for RM 1,000, and pay for transport RM 1,000. The MBA study programme will extend over 18 months. Mr Karim is working now with a salary of RM 3,000 per month. Being a full time student, he will have to forego the income from his present job for the period of study.

 

Work out:  (a) Financial cost, and (b) the Economic Cost of doing MBA

(Hint: Identify implicit costs and explicit costs)


1
Expert's answer
2021-07-08T15:16:59-0400

Mr. Karim is planning to pursue MBA studies. His initial inquiry shows:-

Tuition fees=RM 24000

CostBooks = RM1000

Costtransportation = RM1000

Duration Extended= 18 months

salary= RM3000


"a)\\

 Financial\space Cost= Tuition \space fees +Cost_{books} + Cost_{transportation}\\

 = 24000+1000+1000\\

 = RM\space 26000"

"b) \\

 Economic\space Cost=Financial \space cost+(salary×Duration \space extended)\\ = 26000+(3000×18) \\ = 26000+54000 \\ = RM80000"